### v1

# Linear Complementary Pairs of Quasi-Cyclic and Quasi-Twisted Codes

Abstract
In this paper, we provide a polynomial characterization of linear complementary pairs of quasi-cyclic and quasi-twisted codes of index $2$. We also give several examples of linear complementary pairs of quasi-cyclic and quasi-twisted codes with (almost) optimal security parameters.
Keywords: Quasi-cyclic codes, quasi-twisted codes, linear complementary pairs of codes .
Mathematics subject classification: 94B05, 94B15, 94B60.
1 Introduction
Quasi-cyclic codes are a class of codes that includes cyclic codes. Townsend and Weldon introduced these codes [25] in 1967. It is already known that quasi-cyclic codes are asymptotically good. Beyond classical communication and data storage applications, quasi-cyclic codes are used in modern post-quantum cryptographic systems, such as QC-MDPC code-based protocols, due to their balance of security and compact key sizes. A systematic study of the algebraic structure of quasi-cyclic codes over finite fields and rings has been done in [17, 18, 19, 20]. In 2023, Abdukhalikov et al. [1] investigated one-generator quasi-cyclic codes over finite fields and determined the duals in polynomial representation. Lally and Fitzpatrick [16] gave the polynomial characterization of quasi-cyclic codes by proving that every quasi-cyclic code has a generating set consisting of $\ell$-tuples polynomials with certain properties derived from reduced Gröbner basis. Based on these structural properties, more asymptotic results, minimum distance bounds, and further applications of quasi-cyclic codes were obtained in the literature ([21, 24, 27]). Further, using polynomial characterization, quasi-cyclic codes of index $2$ have been investigated in [2]. Quasi-twisted codes are another generalization of cyclic codes that includes constacyclic and quasi-cyclic codes. Additionally, they are known to be asymptotically good ([10, 26, 15]). Many record-breaking codes with excellent parameters have been constructed within this class (e.g., see [4, 6, 23]).
A pair of linear codes of the same length over a finite field is called a linear complementary pair (LCP) of codes if the codes intersect trivially and the direct sum of the codes is the whole ambient space. LCPs of codes have been extensively studied due to their nice algebraic structure and wide applications in cryptography. Ngo et al. [22] introduced LCPs of codes in 2015 and showed their applications in resistance against physical attacks like side-channel attacks (SCA) and fault injection attacks (FIA). This method is known as Direct Sum Masking (DSM). The security parameters of a given LCP $(C,D)$ over a finite field depend on $\min\{d(C),d(D^{\perp})\}$, where $d(C)$ denotes the minimum Hamming distance of $C$ and $d(D^{\perp})$ denotes the dual distance of $D$ (for details, see [8, 9]). Linear complementary dual (LCD) codes are a special case of an LCP of codes where $D=C^{\perp}$. The security parameters of an LCD code depend only on the minimum distance of the code.
In 2018, Carlet et al. [9] characterized LCPs of quasi-cyclic codes using the Chinese Remainder Theorem (CRT) decomposition of codes and also provided necessary and sufficient conditions for pairs of constacyclic and double circulant codes to be an LCP of codes. Moreover, they proved that if $(C,D)$ is an LCP of constacyclic or 2D cyclic code, then $C$ and $D^{\perp}$ are permutation equivalent. Güneri et al. [12] showed the equivalence of $C$ and $D^{\perp}$ holds for an LCP of $mD$-cyclic codes as well. These results show that finding LCD codes and LCPs of codes is the same in the context of security parameters. However, $C$ and $D^{\perp}$ need not be permutation equivalent for an LCP of quasi-cyclic codes [9, Example 3.3]. In this work, we give a polynomial characterization of LCPs of quasi-cyclic and quasi-twisted codes with index $2$. The rest of the paper is organized as follows.
In Section 2, we recall the basics of LCPs of codes and the decomposition of quasi-cyclic codes. In Section 3, we establish necessary and sufficient conditions for LCPs of quasi-cyclic codes of index $2$. We discuss LCPs of one-generator quasi-cyclic codes in Section 4. We characterize LCPs of quasi-twisted codes in Section 5. In Section 6, we give several examples of LCPs of quasi-cyclic codes.
2 Preliminaries
We denote $\mathbb{F}_{q}$ as a finite field with $q$ elements. A linear code over $\mathbb{F}_{q}$ of length $n$ is a subspace of $\mathbb{F}_{q}^{n}$. We say $C$ is an $[n,k]$ code over $\mathbb{F}_{q}$ code if $C$ is a linear code of length $n$ and $\dim_{\mathbb{F}_{q}}(C)=k$. Let $C$ and $D$ be linear codes of length $n$ over $\mathbb{F}_{q}$. Then $(C,D)$ is called a linear complementary pair (LCP) of codes if they have the trivial intersection and their direct sum is the ambient space $\mathbb{F}_{q}^{n}$. That is,
$C\oplus D=\mathbb{F}_{q}^{n}.$
Note that if $C$ has parameters $[n,k]$, then $D$ must have parameters $[n,n-k]$. Linear complementary dual (LCD) codes are a special class of LCPs of codes where $D$ is the dual code of $C$. Equivalently, $C$ is an LCD code if $C\cap C^{\perp}=\{0\}$.
Let $\lambda\in\mathbb{F}_{q}^{*}$. We define the cyclic $\lambda$-shift operator as
$T_{\lambda}(x_{0},x_{1},\dots,x_{n-1})=(\lambda x_{n-1},x_{0},\dots,x_{n-2}).$
A linear code is said to be a $\lambda$-quasi-twisted code of length $n$ and index $\ell$ (in short, $(\lambda,\ell)$-quasi-twisted) if $T_{\lambda}^{\ell}(c)\in C$ for all $c\in C$. If $\lambda=1$, then $C$ is called a quasi-cyclic code of index $\ell$. When $\ell=1$, then $C$ is called a $\lambda$-constacyclic code, and when $\ell=1$ $\lambda=1$, $C$ is called a cyclic code. If $C$ is a $\lambda$-quasi-twisted code of length $n$ and index $\ell$, then one can assume that $\ell$ divides $n$. Thus, we assume that $n=m\ell$.
Let $R_{\lambda}=\mathbb{F}_{q}[x]/\langle x^{m}-\lambda\rangle$. Note that $\lambda$-constacyclic codes over $\mathbb{F}_{q}$ of length $m$ are ideals of $R_{\lambda}$. Let $C$ be a $\lambda$-quasi-twisted code of length $n=m\ell$ and index $\ell$ over $\mathbb{F}_{q}$. It is well known that $C$ can be identified as an $R_{\lambda}$-submodule of $R_{\lambda}^{\ell}$ (for instance, see [14]). The correspondence is given by
$(c_{0,0},c_{0,1},\dots,c_{0,\ell-1},c_{1,0},\dotsc_{1,\ell-1},\dots,c_{m-1,0},\dots,c_{m-1,\ell-1})\mapsto(c_{0}(x),c_{1}(x),\dots c_{\ell}(x))\in R_{\lambda}^{\ell},$
where $c_{i}(x)=c_{0,i}+c_{1,i}x+\dots+c_{m-1,i}x^{m-1}\in R_{\lambda}$.
Throughout the article, we assume $\gcd(m,q)=1$. We decompose quasi-cyclic codes ($\lambda=1$) into constituent codes (for details, see [13, Chapter 7, Quasi-cyclic codes p.132]). Let $x^{m}-1=f_{1}(x)f_{2}(x)\dots f_{t}(x)$ be the factorization into a product of distinct irreducible polynomials over $\mathbb{F}_{q}$. For each $i$, let $F_{i}:=\mathbb{F}_{q}[x]/\langle f_{i}(x)\rangle$. Let $\xi$ be a primitive $m^{\text{th}}$ root of unity in some extension of $\mathbb{F}_{q}$. Let $\xi^{u_{i}}$ be a root of $f_{i}(x)$. Then $F_{i}=\mathbb{F}_{q}(\xi^{u_{i}})$. By the Chinese Remainder Theorem (CRT), $R=\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle$ can be decomposed as
$R\cong\bigoplus_{i=1}^{t}F_{i}.$
The isomorphism between $R$ and its CRT decomposition is given by
$a(x)\mapsto\left(a(\xi^{u_{1}}),\dots,a(\xi^{u_{t}})\right).$
This isomorphism extends naturally to $R^{\ell}$, which implies that
$R^{\ell}\cong\bigoplus_{i=1}^{t}F_{i}^{\ell}.$
Then, a QC code $C$ of index $\ell$ can be decomposed as
$C\cong\bigoplus_{i=1}^{t}C_{i},$ (2.1)
where each component code is a linear code of length $\ell$ over the base field $F_{i}$. The component codes $C_{i}$ are called the constituents of $C$.
The constituents can be described in terms of the generators of $C$. Namely, if $C$ is an $r$-generator QC code with generators
$\{(a_{1,1}(x),\dots,a_{1,\ell}(x)),\dots,(a_{r,1}(x),\dots,a_{r,\ell}(x))\}\subset R^{\ell},$
then
$C_{i}$ $=\text{Span}_{F_{i}}\{(a_{b,1}(\xi^{u_{i}}),\dots,a_{b,\ell}(\xi^{u_{i}})):1\leq b\leq r\},\text{ for }1\leq i\leq t.$
We have the following characterization of QC LCP of codes from [9, Theorem 3.1].
Theorem 2.1.
Let $C$ and $D$ be $q$-ary QC codes of length $m\ell$ and index $\ell$ with CRT decompositions $C\cong\bigoplus_{i=1}^{t}C_{i}$ and $D\cong\bigoplus_{i=1}^{t}D_{i}$, respectively. Then $(C,D)$ is an LCP of codes if and only if $(C_{i},D_{i})$ is an LCP of codes for all $1\leq i\leq t$.
3 LCPs of quasi-cyclic codes of index $2$
In [16, 2], Lally and Fitzpatrick showed that a quasi-cyclic code of index $\ell$ can be generated by the rows of an upper triangular $\ell\times\ell$ polynomial matrix satisfying certain conditions. We state for $\ell=2$.
Theorem 3.1.
Let $C$ be a quasi-cyclic code of length $2m$ and index $2$ Then $C$ is generated by two elements $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ such that they satisfy the following conditions:
$g_{11}(x)\mid(x^{m}-1)\text{ and }g_{22}(x)\mid(x^{m}-1),$
$\deg g_{12}(x)<\deg g_{22}(x),$ ($\ast$)
$g_{11}(x)g_{22}(x)\mid(x^{m}-1)g_{12}(x).$
Moreover, in this case $\dim C=2m-\deg g_{11}(x)-\deg g_{22}(x)$.
Remark 3.2.
If $\gcd(q,m)=1$, then the condition
$g_{11}(x)g_{22}(x)\mid(x^{m}-1)g_{12}(x)$
in Theorem 3.1 is equivalent to the condition $\gcd(g_{11}(x),g_{22}(x))\mid g_{12}(x)$, since $x^{m}-1$ has no multiple roots.
Let $C$ and $D$ be two quasi-cyclic codes of length $2m$ and index $2$, generated by the rows of the matrices $G$ and $H$, respectively, where
$G=\begin{bmatrix}g_{11}(x)&g_{12}(x)\\ 0&g_{22}(x)\end{bmatrix}\text{ and }H=\begin{bmatrix}f_{11}(x)&f_{12}(x)\\ 0&f_{22}(x)\end{bmatrix}$
and $f_{ij}(x)$, $g_{ij}(x)$ satisfy ()($\ast$)( ∗ ).
Then the constituent codes $C_{i}$ and $D_{i}$ are generated by the rows of the matrices
$G_{i}=\begin{bmatrix}g_{11}(\xi^{u_{i}})&g_{12}(\xi^{u_{i}})\\ 0&g_{22}(\xi^{u_{i}})\end{bmatrix}\text{ and }H_{i}=\begin{bmatrix}f_{11}(\xi^{u_{i}})&f_{12}(\xi^{u_{i}})\\ 0&f_{22}(\xi^{u_{i}})\end{bmatrix},$
respectively.
Let $g(x)=\gcd(g_{11}(x),g_{22}(x))$ and $f(x)=\gcd(f_{11}(x),f_{22}(x))$. Then, by Remark 3.2, $g(x)\mid g_{12}(x)$ and $f(x)\mid f_{12}(x)$. Let $l(x)=(x^{m}-1)/\text{lcm}(g_{11}(x),g_{22}(x))$ and $h(x)=(x^{m}-1)/\text{lcm}(f_{11}(x),f_{22}(x))$. Let $g_{11}(x)=g(x)g_{11}^{\prime}(x),g_{22}(x)=g(x)g_{22}^{\prime}(x)$, and $f_{11}(x)=f(x)f_{11}^{\prime}(x),f_{22}(x)=f(x)f_{22}^{\prime}(x)$. Then
$x^{m}-1=g(x)g_{11}^{\prime}(x)g_{22}^{\prime}(x)l(x)=f(x)f_{11}^{\prime}(x)f_{22}^{\prime}(x)h(x).$
Next, we give a new characterization for a pair of quasi-cyclic codes of index $2$ to be an LCP of codes. We use the notation as above.
Theorem 3.3.
Let $C$ and $D$ be two quasi-cyclic codes of length $2m$ and index $2$ generated by $(g_{11}(x),g_{12}(x))$, $(0,g_{22}(x))$ and $(f_{11}(x),f_{12}(x))$, $(0,f_{22}(x))$, respectively, satisfying ()($\ast$)( ∗ ). Then $(C,D)$ is an LCP of codes if and only if all of the following conditions are true:
  1. (I)
    $\gcd(f_{11}(x),g_{11}(x))=1$.
  2. (II)
    $g(x)=\frac{x^{m}-1}{\mathrm{lcm}(f_{11}(x),f_{22}(x))}$.
  3. (III)
    $f(x)=\frac{x^{m}-1}{\mathrm{lcm}(g_{11}(x),g_{22}(x))}$.
  4. (IV)
    $\gcd(g_{22}^{\prime}(x),f_{22}^{\prime}(x),g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))=1$.
We provide the proof of Theorem 3.3 through the following sequence of lemmata.
Lemma 3.4.
If $(C,D)$ is an LCP of codes, then (I) holds.
Proof.
Assume that $\gcd(f_{11}(x),g_{11}(x))\neq 1$. This implies that there exists $f_{i}(x)$ such that $f_{i}(x)\mid g_{11}(x)$ and $f_{i}(x)\mid f_{11}(x)$.
Then $g_{11}(\xi^{u_{i}})=f_{11}(\xi^{u_{i}})=0$, and $C_{i}$ and $D_{i}$ are generated by the rows of the matrices
$G_{i}=\begin{bmatrix}g_{11}(\xi^{u_{i}})&g_{12}(\xi^{u_{i}})\\ 0&g_{22}(\xi^{u_{i}})\end{bmatrix}=\begin{bmatrix}0&g_{12}(\xi^{u_{i}})\\ 0&g_{22}(\xi^{u_{i}})\end{bmatrix}\text{ and }H_{i}=\begin{bmatrix}f_{11}(\xi^{u_{i}})&f_{12}(\xi^{u_{i}})\\ 0&f_{22}(\xi^{u_{i}})\end{bmatrix}=\begin{bmatrix}0&f_{12}(\xi^{u_{i}})\\ 0&f_{22}(\xi^{u_{i}})\end{bmatrix},$
respectively. This implies that $\dim(C_{i}+D_{i})\neq 2$. Therefore, $(C_{i},D_{i})$ is not an LCP of codes. ∎
Lemma 3.5.
If $(C,D)$ is an LCP of codes, then (II) holds.
Proof.
By Lemma 3.4, $(I)$ holds. Assume $g(x)\neq\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$. This implies there exist irreducible $f_{i}(x)$ or $f_{j}(x)$ such that
$(i)$ $f_{i}(x)\mid g(x)$ and $f_{i}(x)\nmid\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$,
$(ii)$ $f_{j}(x)\nmid g(x)$ and $f_{j}(x)\mid\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$.
$(i)$ If $f_{i}(x)\mid g(x)$ and $f_{i}(x)\nmid\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$, then $f_{i}(x)\mid g_{11}(x),g_{22}(x),g_{12}(x)$ and $f_{i}(x)\mid\text{lcm}(f_{11}(x),f_{22}(x))$. This implies that $f_{i}(x)\mid f_{22}(x)$ since $\gcd(g_{11}(x),f_{11}(x))=1$. Thus, $g_{11}(\xi^{u_{i}})=g_{22}(\xi^{u_{i}})=g_{12}(\xi^{u_{i}})=f_{22}(\xi^{u_{i}})=0$ and $C_{i}$ and $D_{i}$ are generated by the rows of the matrices,
$G_{i}=\begin{bmatrix}g_{11}(\xi^{u_{i}})&g_{12}(\xi^{u_{i}})\\ 0&g_{22}(\xi^{u_{i}})\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\end{bmatrix}\text{ and }H_{i}=\begin{bmatrix}f_{11}(\xi^{u_{i}})&f_{12}(\xi^{u_{i}})\\ 0&f_{22}(\xi^{u_{i}})\end{bmatrix}=\begin{bmatrix}f_{11}(\xi^{u_{i}})&f_{12}(\xi^{u_{i}})\\ 0&0\end{bmatrix},$
respectively. This implies that $\dim(C_{i})+\dim(D_{i})\neq 2$. Therefore, $(C_{i},D_{i})$ is not an LCP of codes.
$(ii)$ If $f_{j}(x)\nmid g(x)$ and $f_{j}(x)\mid\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$, then $f_{j}(x)\nmid g_{11}(x)$ or $f_{j}(x)\nmid g_{22}(x)$, and $f_{j}(x)\nmid\text{lcm}(f_{11}(x),f_{22}(x))$ (equivalently, $f_{j}(x)\nmid f_{11}(x),f_{22}(x)$). This implies that $g_{11}(\xi^{u_{j}})\neq 0$ or $g_{22}(\xi^{u_{j}})\neq 0$, and $f_{11}(\xi^{u_{j}})\neq 0\neq f_{22}(\xi^{u_{j}})$. Then $C_{j}$ and $D_{j}$ are generated by the rows of the matrix
$G_{j}=\begin{bmatrix}g_{11}(\xi^{u_{j}})&g_{12}(\xi^{u_{j}})\\ 0&g_{22}(\xi^{u_{j}})\end{bmatrix}\text{ and }H_{j}=\begin{bmatrix}f_{11}(\xi^{u_{j}})&f_{12}(\xi^{u_{j}})\\ 0&f_{22}(\xi^{u_{j}})\end{bmatrix},$
respectively. This implies that $\dim(C_{j})\geq 1$ and $\dim(D_{j})=2$. Thus $C_{j}\cap D_{j}\neq\{0\}$. Therefore, $(C_{j},D_{j})$ is not an LCP of codes. ∎
Lemma 3.6.
If $(C,D)$ is an LCP of codes, then $(III)$ holds.
Proof.
The proof is similar to that of Lemma 3.5
Lemma 3.7.
If $(C,D)$ is an LCP of codes, then $(IV)$ holds.
Proof.
Assume $\gcd(g_{22}^{\prime}(x),f_{22}^{\prime}(x),g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))\neq 1$. Then there exists $f_{i}(x)$ such that $f_{i}(x)\mid g_{22}^{\prime}(x),f_{22}^{\prime}(x)$ and $f_{i}(x)\mid(g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))$. This implies that $g_{22}^{\prime}(\xi^{u_{i}})=f_{22}^{\prime}(\xi^{u_{i}})=g_{11}(\xi^{u_{i}})f_{12}(\xi^{u_{i}})-g_{12}(\xi^{u_{i}})f_{11}(\xi^{u_{i}})=0$ and $C_{i}$ and $D_{i}$ are generated by the rows of the matrices
$G_{i}=\begin{bmatrix}g_{11}(\xi^{u_{i}})&g_{12}(\xi^{u_{i}})\\ 0&g_{22}(\xi^{u_{i}})\end{bmatrix}=\begin{bmatrix}g_{11}(\xi^{u_{i}})&g_{12}(\xi^{u_{i}})\\ 0&0\end{bmatrix}$
[…]
respectively. Then $C_{i}\cap D_{i}=\{0\}$ if and only if the determinant of the matrix $A=\begin{bmatrix}g_{11}(\xi^{u_{i}})&g_{12}(\xi^{u_{i}})\\ f_{11}(\xi^{u_{i}})&f_{12}(\xi^{u_{i}})\end{bmatrix}$ is non-zero, but $g_{11}(\xi^{u_{i}})f_{12}(\xi^{u_{i}})-g_{12}(\xi^{u_{i}})f_{11}(\xi^{u_{i}})=0$ implies that $det(A)=0$. Thus $C_{i}\cap D_{i}\neq\{0\}$. Therefore, $(C_{i},D_{i})$ is not an LCP of codes. ∎
Lemma 3.8.
If $(I),(II),(III),(IV)$ hold, then $(C,D)$ is an LCP of codes, that is $C_{i}\cap D_{i}=\{0\}$ and $\dim(C_{i})+\dim(D_{i})=2$ for all $1\leq i\leq t$.
Proof.
Let $l(x)=(x^{m}-1)/\text{lcm}(g_{11}(x),g_{22}(x))$ and $h(x)=(x^{m}-1)/\text{lcm}(f_{11}(x),f_{22}(x))$. By $(II)$ and $(III)$, we get
$x^{m}-1=g(x)g_{11}^{\prime}(x)g_{22}^{\prime}(x)f(x)=g(x)f_{11}^{\prime}(x)f_{22}^{\prime}(x)f(x).$
For each $i$, the irreducible polynomial $f_{i}$ divides exactly one of the factors $g(x)$, $g_{11}^{\prime}(x)$, $g_{22}^{\prime}(x)$, $f(x)$ of $x^{m}-1$, which leads to the following four cases.
$(i)$ $f_{i}(x)\mid g(x)$. Then $f_{i}(x)\mid g_{11}(x),g_{22}(x),g_{12}(x)$ and $f_{i}(x)\nmid f_{11}^{\prime}(x),f_{22}^{\prime}(x),f(x)$. It follows that
$G_{i}=\begin{bmatrix}0&0\\ 0&0\end{bmatrix},\text{ and }H_{i}=\begin{bmatrix}f_{11}(\xi^{u_{i}})&f_{12}(\xi^{u_{i}})\\ 0&f_{22}(\xi^{u_{i}})\end{bmatrix},$
where $f_{11}(\xi^{u_{i}})\neq 0\neq f_{22}(\xi^{u_{i}})$. It can then be readily checked that $C_{i}\cap D_{i}=\{0\}$ and $\dim(C_{i})+\dim(D_{i})=2$.
$(ii)$ $f_{i}(x)\mid g_{11}^{\prime}(x)$. This implies that $f_{i}(x)\nmid g(x),g_{22}^{\prime}(x),f(x)$ and then either $f_{i}(x)\mid f_{11}^{\prime}(x)$ or $f_{i}(x)\mid f_{22}^{\prime}(x)$. By $(I)$, $f_{i}(x)\nmid f_{11}(x)$ and thus $f_{i}(x)\mid f_{22}^{\prime}(x)$. It follows that
$G_{i}=\begin{bmatrix}0&g_{12}(\xi^{u_{i}})\\ 0&g_{22}(\xi^{u_{i}})\end{bmatrix},\text{ and }H_{i}=\begin{bmatrix}f_{11}(\xi^{u_{i}})&f_{12}(\xi^{u_{i}})\\ 0&0\end{bmatrix},$
where $g_{22}(\xi^{u_{i}})\neq 0\neq f_{11}(\xi^{u_{i}})$. It can then be readily checked that $C_{i}\cap D_{i}=\{0\}$ and $\dim(C_{i})+\dim(D_{i})=2$.
$(iii)$ $f_{i}(x)\mid g_{22}^{\prime}(x)$. This implies that $f_{i}(x)\nmid g(x),g_{11}^{\prime}(x),f(x)$ and thus either $f_{i}(x)\mid f_{11}^{\prime}(x)$ or $f_{i}(x)\mid f_{22}^{\prime}(x)$.
1. If $f_{i}(x)\mid f_{11}^{\prime}(x)$ then $f_{i}(x)\nmid f_{22}^{\prime}(x)$ and thus
$G_{i}=\begin{bmatrix}g_{11}(\xi^{u_{i}})&g_{12}(\xi^{u_{i}})\\ 0&0\end{bmatrix}\text{ and }H_{i}=\begin{bmatrix}0&f_{12}(\xi^{u_{i}})\\ 0&f_{22}(\xi^{u_{i}})\end{bmatrix},$
where $g_{11}(\xi^{u_{i}})\neq 0\neq f_{22}(\xi^{u_{i}})$. It can then be readily checked that $C_{i}\cap D_{i}=\{0\}$ and $\dim(C_{i})+\dim(D_{i})=2$.
2. If $f_{i}(x)\mid f_{22}^{\prime}(x)$ then $f_{i}(x)\nmid f_{11}^{\prime}(x)$ and thus
$G_{i}=\begin{bmatrix}g_{11}(\xi^{u_{i}})&g_{12}(\xi^{u_{i}})\\ 0&0\end{bmatrix}\text{ and }H_{i}=\begin{bmatrix}f_{11}(\xi^{u_{i}})&f_{12}(\xi^{u_{i}})\\ 0&0\end{bmatrix}.$
By $(IV)$, $f_{i}(x)\nmid(g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))$ (since $f_{i}(x)\mid g_{22}^{\prime}(x),f_{22}^{\prime}(x)$). This implies that $g_{11}(\xi^{u_{i}})f_{12}(\xi^{u_{i}})-g_{12}(\xi^{u_{i}})f_{11}(\xi^{u_{i}})\neq 0$. Therefore, $C_{i}\cap D_{i}=\{0\}$ and $\dim(C_{i})+\dim(D_{i})=2$.
$(iv)$ $f_{i}(x)\mid f(x)$. This implies that, $f_{i}(x)\nmid g(x),g_{11}^{\prime}(x),g_{22}^{\prime}(x)$ and $f_{i}(x)\mid f_{11}(x),f_{22}(x),f_{12}(x)$. Thus
$G_{i}=\begin{bmatrix}g_{11}(\xi^{u_{i}})&g_{12}(\xi^{u_{i}})\\ 0&g_{22}(\xi^{u_{i}})\end{bmatrix},\text{ and }H_{i}=\begin{bmatrix}0&0\\ 0&0\end{bmatrix},$
where $g_{11}(\xi^{u_{i}})\neq 0\neq g_{22}(\xi^{u_{i}})$. It can then be readily checked that $C_{i}\cap D_{i}=\{0\}$ and $\dim(C_{i})+\dim(D_{i})=2$.
Proof of Theorem 3.3.
We recall that we want to prove that $(C,D)$ is an LCP of codes if and only if the conditions (I), (II), (III), and (IV) hold. The “if” direction follows from Lemmata 3.8. The “only if” direction follows from Lemmata 3.4, 3.5, 3.6, and 3.7.
4 LCPs of one-generator quasi-cyclic codes of index 2
In this section, we consider one-generator quasi-cyclic codes of index $2$. The following lemma from [3, Lemma 2.4] shows how one-generator quasi-cyclic codes are generated by two generators satisfying ()($\ast$)( ∗ ).
Lemma 4.1.
Let $\gcd(q,m)=1$ and let $C$ be a quasi-cyclic code generated by one element $(g_{11}(x),g_{12}(x))$, where $g_{11}(x)\mid(x^{m}-1)$. Let $g(x)=\gcd(g_{11}(x),g_{12}(x))$, $g_{11}(x)=g(x)g_{11}^{\prime}(x)$, $g_{12}(x)=g(x)g_{12}^{\prime}(x)$. Let
$g_{22}(x)=\dfrac{x^{m}-1}{g_{11}^{\prime}(x)}.$
Then the code $C$ is generated by two elements $(g_{11}(x),g_{12}(x)\mod{g_{22}(x)})$ and $(0,g_{22}(x))$ satisfying ()($\ast$)( ∗ ) with $\gcd(g_{11}(x),g_{22}(x))=g(x)$.
From the above Lemma 4.1, if $C$ is one-generator quasi-cyclic code of index $2$, then we can assume that $C$ is generated by two elements $(g_{11}(x),\tilde{g}_{12}(x))$ and $(0,g_{22}(x))$ satisfying ()($\ast$)( ∗ ), where $\tilde{g}_{12}(x)=g_{12}(x)\mod{g_{22}(x)}$. Furthermore, also by Lemma 4.1 we have that $\gcd(g_{11}(x),g_{22}(x))=g(x)$. In addition, $(0,g_{22}(x))\in C$ (see proof of [3, Lemma 2.4]). It follows that the code $C$ generated by the element $(g_{11}(x),g_{12}(x))$ is the same as the code generated by the element $(g_{11}(x),\tilde{g}_{12}(x))$. Thus, without loss of generality, from now on, we will write $g_{12}(x)$ instead of $\tilde{g}_{12}(x)$. We have
$\text{lcm}(g_{11}(x),g_{22}(x))=\dfrac{g_{11}(x)\cdot g_{22}(x)}{\gcd(g_{11}(x),g_{22}(x))}=\dfrac{g(x)\cdot g_{11}^{\prime}(x)\cdot(x^{m}-1)}{g(x)\cdot g_{11}^{\prime}(x)}=x^{m}-1.$ (4.1)
Lemma 4.2.
Let $g_{11}(x)$, $g_{12}(x)$, $f_{11}(x)$, and $f_{12}(x)$ in $\mathbb{F}_{q}[x]$ be such that $g_{11}(x)$ and $f_{11}(x)$ divide $x^{m}-1$. Let $\gcd(g_{11}(x),g_{12}(x))=1$, $\gcd(f_{11}(x),f_{12}(x))=1$ and $\gcd(f_{11}(x),g_{11}(x))=1$. Then the following are equivalent
  1. (1)
    $\gcd\left(\frac{x^{m}-1}{f_{11}(x)},\frac{x^{m}-1}{g_{11}(x)},g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x)\right)=1$
  2. (2)
    $\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))=1$
Proof.
$(2)$ implies $(1)$ trivially. Let $\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))\neq 1$. Then there exists irreducible $f_{i}(x)$ such that $f_{i}(x)$ divides both $x^{m}-1$ and $g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x)$. Assume that $f_{i}(x)|g_{11}(x)$. Then $f_{i}(x)|g_{12}(x)f_{11}(x)$, which implies that $f_{i}(x)|g_{12}(x)$ or $f_{i}(x)|f_{11}(x)$, which is not possible. Thus $f_{i}(x)\nmid g_{11}(x)$. Similarly, $f_{i}(x)\nmid f_{11}(x)$. Hence, $\gcd\left(\frac{x^{m}-1}{f_{11}(x)},\frac{x^{m}-1}{g_{11}(x)},g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x)\right)\neq 1$. This completes the proof. ∎
Theorem 4.3.
Let $C$ and $D$ be one-generator quasi-cyclic codes generated by $(g_{11}(x),g_{12}(x))$, and $(f_{11}(x),f_{12}(x))$, respectively, where $g_{11}(x),f_{11}(x)\mid(x^{m}-1)$. Then $(C,D)$ is an LCP of codes if and only if all the following conditions hold:
  1. (A)
    $\gcd(g_{11}(x),g_{12}(x))=1$.
  2. (B)
    $\gcd(f_{11}(x),f_{12}(x))=1$.
  3. (C)
    $\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x)=1$.
Proof.
Let $g(x)=\gcd(g_{11}(x),g_{12}(x))$, $g_{11}(x)=g(x)g_{11}^{\prime}(x)$, $f(x)=\gcd(f_{11}(x),f_{12}(x))$ and $f_{11}(x)=f(x)f_{11}^{\prime}(x)$. Then by Lemma 4.1, we can assume that $C$ is generated by two elements $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$, and $D$ is generated by two elements $(f_{11}(x),f_{12}(x))$ and $(0,f_{22}(x))$, where
$g_{22}(x)=\frac{x^{m}-1}{g_{11}^{\prime}(x)}\text{ and }f_{22}(x)=\frac{x^{m}-1}{f_{11}^{\prime}(x)}.$
It is easy to see the generating elements for $C$ and $D$ satisfy the conditions ()($\ast$)( ∗ ) and $x^{m}-1=\text{lcm}(g_{11}(x),g_{22}(x))=\text{lcm}(f_{11}(x),f_{22}(x))$ (see Eq. 4.1). By Theorem 3.3, $(C,D)$ is an LCP of codes if and only if
  1. (i)
    $\gcd(f_{11}(x),g_{11}(x))=1$.
  2. (ii)
    $g(x)=1$.
  3. (iii)
    $f(x)=1$.
  4. (iv)
    $\gcd\left(\frac{x^{m}-1}{g_{11}(x)},\frac{x^{m}-1}{f_{11}(x)},g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x)\right)=1$.
We show that $(i),(ii),(iii),(iv)$ are equivalent to $(A),(B),(C)$. First, assume that $(i)$, $(ii)$, $(iii)$, and $(iv)$ hold. Then, by Lemma 4.2 $(C)$ holds. Also, $(ii),(iii)$ are the same as $(A)$ and $(B)$.
Conversely, suppose $(A),(B),(C)$ hold. Then $(ii)$, $(iii)$ and $(iv)$ hold trivially. If $\gcd(g_{11}(x),f_{11}(x))\neq 1$, then $\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))\neq 1$. Hence $(C)$ implies $(i)$. This completes the proof. ∎
In [5], the authors called a code $C$ maximal if $\dim(C)=m$. Recall that if $C$ is a one-generator quasi-cyclic code of length $2m$ and index $2$, generated by $(g_{11}(x),g_{12}(x))$ with $g_{11}(x)\mid(x^{m}-1)$, then $\dim(C)=m-\deg\ g(x)$, where $g(x)=\gcd(g_{11}(x),g_{12}(x))$.
Corollary 4.4.
[5, Proposition 3.6] Let $C=\langle(g_{11}(x),g_{12}(x))\rangle$ and $D=\langle(f_{11}(x),f_{12}(x))\rangle$ be maximal one-generator quasi-cyclic codes of index $2$, where $g_{11}(x),f_{11}(x)\mid(x^{m}-1)$. Then $(C,D)$ is an LCP of codes if and only if $\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))=1$.
Let $R=\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle$. Double circulant (DC) codes are one-generator quasi-cyclic codes of index $2$, generated by $\langle(1,a(x))\rangle\in R^{2}$. In the next result, we show that Theorem 4.3 generalizes [9, Proposition 3.2].
Corollary 4.5.
Let $C=\langle(1,a(x))\rangle$ and $D=\langle(1,b(x))\rangle$ be DC codes in $R^{2}$. Then $(C,D)$ is an LCP of codes if and only if $\gcd(a(x)-b(x),x^{m}-1)=1$.
Proof.
Take $g_{11}(x)=1,g_{12}(x)=a(x),f_{11}(x)=1,f_{12}(x)=b(x)$. Then, the proof follows from Theorem 4.3. ∎
5 LCPs of quasi-twisted codes of index 2
In this section, we characterize the LCP of quasi-twisted codes of index $2$. First, we describe the structure of quasi-twisted codes of index $2$ analogous to that of quasi-cyclic codes.
Theorem 5.1.
Let $C$ be a $\lambda$-quasi-twisted code of length $2m$ and index $2$ Then $C$ is generated by two elements $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ such that they satisfy the following conditions:
$g_{11}(x)\mid(x^{m}-\lambda)\text{ and }g_{22}(x)\mid(x^{m}-\lambda),$
[…]
$g_{11}(x)g_{22}(x)\mid(x^{m}-\lambda)g_{12}(x).$
Moreover, in this case $\dim C=2m-\deg g_{11}(x)-\deg g_{22}(x)$.
Proof.
In [16] quasi-cyclic codes (i.e., quasi-twisted codes with $\lambda=1$) were studied using Gröbner bases of modules. We note that the results of [16] are valid for any $\lambda$ (all proofs are valid if we take $x^{m}-\lambda$ instead of $x^{m}-1$). Hence, by [16] we can assume that $C$ is generated by two elements $g_{1}=\big{(}g_{11}(x),g_{12}(x)\big{)}$ and $g_{2}=\big{(}0,g_{22}(x)\big{)}$, satisfying the conditions $g_{11}(x)\mid(x^{m}-\lambda)$, $g_{22}(x)\mid(x^{m}-\lambda)$, $\deg g_{12}(x)<\deg g_{22}(x)$, and $\dim C=2m-\deg g_{11}(x)-\deg g_{22}(x)$. The existence of such generators is equivalent [16] to the existence of a $2\times 2$ polynomial matrix $(a_{ij})$ such that
$\left(\begin{array}[]{cc}a_{11}&a_{12}\\ a_{21}&a_{22}\end{array}\right)\left(\begin{array}[]{cc}g_{11}&g_{12}\\ 0&g_{22}\end{array}\right)=\left(\begin{array}[]{cc}x^{m}-\lambda&0\\ 0&x^{m}-\lambda\end{array}\right).$
Then $a_{21}=0$, $a_{11}=\frac{x^{m}-\lambda}{g_{11}(x)}$, $a_{22}=\frac{x^{m}-\lambda}{g_{22}(x)}$, and $\frac{x^{m}-\lambda}{g_{11}(x)}g_{12}+a_{12}g_{22}=0$, which implies $g_{11}(x)g_{22}(x)\mid(x^{m}-\lambda)g_{12}(x)$. ∎
Remark 5.2.
If $\gcd(q,m)=1$, then the condition
$g_{11}(x)g_{22}(x)\mid(x^{m}-\lambda)g_{12}(x)$
in Theorem 5.1 is equivalent to the condition $\gcd(g_{11}(x),g_{22}(x))\mid g_{12}(x)$, since $x^{m}-\lambda$ has no multiple roots.
Recall that a $\lambda$-quasi-twisted code of length $\ell m$ and index $\ell$ is an $R_{\lambda}$-submodule of $R_{\lambda}^{\ell}$. Let $x^{m}-\lambda=f_{1}(x)f_{2}(x)\dots f_{s}(x)$ be its irreducible factorization over $\mathbb{F}_{q}$. Then, we can decompose quasi-twisted codes into their constituent codes analogous to the quasi-cyclic case (for details, see [14]). A quasi-twisted code $C$ of index $\ell$ can be decomposed as
$C\cong\bigoplus_{i=1}^{s}C_{i}.$
In the next theorem, we state a characterization of an LCP of quasi-twisted codes. The proof is similar to that for an LCP of quasi-cyclic codes given in [9]. We omit the proof.
Theorem 5.3.
Let $C$ and $D$ be $q$-ary $\lambda$-QT codes of length $m\ell$ and index $\ell$ with CRT decompositions $C\cong\bigoplus_{i=1}^{s}C_{i}$ and $D\cong\bigoplus_{i=1}^{s}D_{i}$, respectively. Then $(C,D)$ is an LCP of quasi-twisted codes if and only if $(C_{i},D_{i})$ is an LCP of codes for all $1\leq i\leq s$.
Let $g(x)=\gcd(g_{11}(x),g_{22}(x))$ and $f(x)=\gcd(f_{11}(x),f_{22}(x))$. Then, by Remark 5.2, $g(x)\mid g_{12}(x)$ and $f(x)\mid f_{12}(x)$. Let $g_{11}(x)=g(x)g_{11}^{\prime}(x),g_{22}=g(x)g_{22}^{\prime}(x)$, and $f_{11}(x)=f(x)f_{11}^{\prime}(x),f_{22}=f(x)f_{22}^{\prime}(x)$. Next, we provide another characterization of two-generator and one-generator LCPs of quasi-twisted codes of index $2$.
Theorem 5.4.
[…]
  1. (I)
    $\gcd(f_{11}(x),g_{11}(x))=1$.
  2. (II)
    $g(x)=\frac{x^{m}-\lambda}{\mathrm{lcm}(f_{11}(x),f_{22}(x))}$.
  3. (III)
    $f(x)=\frac{x^{m}-\lambda}{\mathrm{lcm}(g_{11}(x),g_{22}(x))}$.
  4. (IV)
    $\gcd(g_{22}^{\prime}(x),f_{22}^{\prime}(x),g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))=1$.
Proof.
The proof is similar to that for Theorem 3.3. ∎
Theorem 5.5.
Let $C$ and $D$ be one-generator $\lambda$-quasi-twisted codes generated by $(g_{11}(x),g_{12}(x))$, and $(f_{11}(x),f_{12}(x))$ respectively, where $g_{11}(x),f_{11}(x)\mid(x^{m}-\lambda)$. Then $(C,D)$ is an LCP of codes if and only if
  1. (A)
    $\gcd(g_{11}(x),g_{12}(x))=1$.
  2. (B)
    $\gcd(f_{11}(x),f_{12}(x))=1$.
  3. (C)
    $\gcd(x^{m}-\lambda,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x)=1$.
Proof.
The proof is similar to that for Theorem 4.3. ∎
6 Examples of LCPs of codes with good security parameters
In this section, we present several examples of LCPs of codes with security parameters. The parameters of a linear code are denoted by $[n,k,d]$, where $n$ is the length, $k$ is the dimension, and $d$ is the minimum Hamming distance of the code. We denote the security parameter of an LCP of codes $(C,D)$ by $d_{LCP}=\min\{d(C),d(D^{\perp})\}$ for given values of $n,k$, and $d_{BKLC}$ denotes the minimum distance of the best known linear code for given $n,k$. All computations are done in MAGMA [7]. For the values of $d_{BKLC}$, we refer to Grassl’s code table [11].
Example 6.1.
Let $q=5$ and $m=8$. Then $x^{m}-1=(x+1)(x+2)(x+3)(x+4)(x^{2}+2)(x^{2}+3)$ in $\mathbb{F}_{5}[x]$. Suppose $C$ is generated by
$\begin{split}g_{11}(x)&=x^{2}+3x+2,\\ g_{12}(x)&=x^{5}+4x^{4}+4x^{3}+x^{2}+4x+4,\\ g_{22}(x)&=x^{6}+x^{5}+3x^{4}+x^{2}+x+3\end{split}$
and $D$ is generated by
$\begin{split}f_{11}(x)&=x+3,\\ f_{12}(x)&=x^{3}+2x^{2}+x+2,\\ f_{22}(x)&=x^{7}+3x^{6}+4x^{5}+2x^{4}+x^{3}+3x^{2}+4x+2.\end{split}$
Then $(C,D)$ is an LCP of quasi-cyclic codes with $n=16$, $k=8$ and $d_{LCP}=6$. Note that for $n=16$, $k=8$, we have $d_{BKLC}=7$.
Example 6.2.
Let $q=3$ and $m=4$. Suppose $C$ is generated by
$\begin{split}g_{11}(x)&=x+1,\\ g_{12}(x)&=x^{2}+x+1,\\ g_{22}(x)&=x^{3}+2x^{2}+x+2\end{split}$
and $D$ is generated by
$\begin{split}f_{11}(x)&=x+2,\\ f_{12}(x)&=x^{2}+x+2,\\ f_{22}(x)&=x^{3}+x^{2}+x+1.\end{split}$
Then $(C,D)$ is an LCP of ternary quasi-cyclic codes with $n=8$, $k=4$ and $d_{LCP}=4$. This security parameter is equal to the minimum distance $d_{BKLC}$ of the best known linear code.
Example 6.3.
Let $q=3$ and $m=5$. Suppose $C$ is generated by
$\begin{split}g_{11}(x)&=x+2,\\ g_{12}(x)&=x^{2}+2x+2,\\ g_{22}(x)&=x^{4}+x^{3}+x^{2}+x+1\end{split}$
and $D$ is generated by
$\begin{split}f_{11}(x)&=1,\\ f_{12}(x)&=x^{3}+2x^{2}+2x+1,\\ f_{22}(x)&=x^{5}-1.\end{split}$
Then $(C,D)$ is an LCP of ternary quasi-cyclic codes with $n=10$, $k=5$ and $d_{LCP}=5=d_{BKLC}$. This security parameter improves the security parameter obtained in [5, Table 3] with $n=10$ $k=5$.
In Table 1, we list some examples of LCPs of quasi-cyclic codes $(C,D)$ over $\mathbb{F}_{4}=\mathbb{F}_{2}[w]$, where the code $C$ is generated by $(g_{11}(x),g_{12}(x)),(0,g_{22}(x))$ and the code $D$ is generated by $(f_{11}(x),f_{12}(x)),(0,f_{22}(x))$. We compare the security parameter $d_{LCP}$ with $d_{BKLC}$.
$n$ $k$ $g_{11}(x)$ $g_{12}(x)$ $g_{22}(x)$ $f_{11}(x)$ $f_{12}(x)$ $f_{22}(x)$ $d_{LCP}$ $d_{BKLC}$
$6$ $3$ $x+w$ $x+1+w^{2}$ $x^{2}+wx$ $+w^{2}$ $1$ $wx^{3}+wx^{2}$ $+w^{2}x+1$ $x^{3}-1$ $3$ $4$
$10$ $5$ $x+1$ $x^{3}+wx^{2}$ $+wx$ $x^{4}+x^{3}$ $+x^{2}+x$ $+1$ $1$ $x^{4}+wx^{3}$ $+x+w$ $x^{5}-1$ $5$ $5$
$14$ $9$ $x+1$ $x^{3}+w^{2}x$ $+w$ $x^{4}+x^{3}$ $+x^{2}+1$ $x^{3}+x^{2}$ $+1$ $x^{3}+x^{2}$ $+1$ $x^{6}+x^{5}$ $+x^{4}+x^{3}$ $+x^{2}+x+1$ $4$ $4$
Table 1: LCPs of quasi-cyclic codes over $\mathbb{F}_{4}$
Next, we provide some examples of LCPs of quasi-twisted codes with good security parameters. In the following example, we get the optimal security parameter over $\mathbb{F}_{5}$ with $n=22$ and $k=11$.
Example 6.4.
Let $q=5,$ $m=11$ and $\lambda=2$. Suppose $C$ is a $\lambda$-quasi-twisted code generated by
$\begin{split}g_{11}(x)&=x+2,\\ g_{12}(x)&=x^{6}+4x^{5}+2x^{3}+3x^{2}+x+3,\\ g_{22}(x)&=x^{10}+3x^{9}+4x^{8}+2x^{7}+x^{6}+3x^{5}+4x^{4}+2x^{3}+x^{2}+3x+4\end{split}$
and $D$ is an another $\lambda$-quasi-twisted code generated by
$\begin{split}f_{11}(x)&=1,\\ f_{12}(x)&=x^{9}+2x^{8}+x^{7}+2x^{6}+4x^{5}+2x^{4}+x^{2}+2x+1,\\ f_{22}(x)&=x^{11}-2.\end{split}$
Then $(C,D)$ is an LCP of $\lambda$-quasi-twisted quinary codes with $n=22$, $k=11$ and $d_{LCP}=8$. This security parameter $d_{LCP}$ is the same as the minimum distance of the best known linear codes $d_{BKLC}=8$.
Example 6.5.
Let $q=9,$ $m=10$ and $\lambda=w$, where $\mathbb{F}_{9}=\mathbb{F}_{3}[w]$ and $w^{2}+2w+2=0$. Suppose $C$ is a $\lambda$-quasi-twisted code generated by
$\begin{split}g_{11}(x)&=x^{2}+2x+w,\\ g_{12}(x)&=x^{5}+w^{2}x^{4}+w^{5}x^{3}+w^{6}x^{2}+w^{2}x+w^{3},\\ g_{22}(x)&=x^{8}+x^{7}+w^{3}x^{6}+w^{2}x^{5}+w^{7}x^{4}+w^{3}x^{3}+w^{5}x^{2}+w^{3}x+2\end{split}$
and $D$ is an another $\lambda$-quasi-twisted code generated by
$\begin{split}f_{11}(x)&=x^{2}+w^{7}x+w,\\ f_{12}(x)&=x^{7}+w^{7}x^{6}+wx^{5}+w^{6}x^{4}+x^{3}+w^{7}x^{2}+w^{3}x,\\ f_{22}(x)&=x^{8}+w^{3}x^{7}+w^{7}x^{6}+w^{7}x^{5}+wx^{4}+x^{3}+wx^{2}+w^{6}x+2.\end{split}$
Then $(C,D)$ is an LCP of $\lambda$-quasi-twisted codes over $\mathbb{F}_{9}$ with $n=20$, $k=10$ and $d_{LCP}=8$. The minimum distance of the best known linear codes is $d_{BKLC}=10$.
Example 6.6.
Let $q=9,$ $m=10$ and $\lambda=w$, where $\mathbb{F}_{9}=\mathbb{F}_{3}[w]$ and $w^{2}+2w+2=0$. Suppose $C$ is a $\lambda$-quasi-twisted code generated by
$\begin{split}g_{11}(x)&=1,\\ g_{12}(x)&=x^{5}+x^{4}+wx^{3}+w^{5}x^{2}+w^{2},\\ g_{22}(x)&=x^{6}+w^{3}x^{5}+2x^{4}+wx^{3}+w^{5}x^{2}+w^{5}x+w^{3}\end{split}$
and $D$ is an another $\lambda$-quasi-twisted code generated by
$\begin{split}f_{11}(x)&=x^{4}+w^{7}x^{3}+w^{5}x^{2}+x+w^{2},\\ f_{12}(x)&=w^{6}x^{8}+w^{3}x^{7}+w^{7}x^{6}+w^{6}x^{5}+w^{7}x^{4}+w^{3}x^{3}+x^{2}+w^{5}x,\\ f_{22}(x)&=x^{10}+w^{5}.\end{split}$
Then $(C,D)$ is an LCP of $\lambda$-quasi-twisted codes over $\mathbb{F}_{9}$, where $C$ is a $[20,14]$ code and $d_{LCP}=5$. The minimum distance of the best known linear codes with $n=20$ and $k=14$ is $d_{BKLC}=6$.
Conclusion
Linear complementary pairs (LCP) of codes generalize linear complementary dual (LCD) codes and have applications in resistance against side-channel and fault-injection attacks. In this work, we presented a polynomial characterization of LCPs of quasi-cyclic and quasi-twisted codes of index $2$. As a consequence, we provided necessary and sufficient conditions for LCPs of one-generator quasi-cyclic and quasi-twisted codes. In addition, we provided several examples of LCPs of quasi-cyclic and quasi-twisted codes with (almost) optimal security parameters.
Acknowledgment
K. Abdukhalikov was supported by UAEU grants G00004233 and G00004614. D. Ho was supported by the Tromsø Research Foundation (project “Pure Mathematics in Norway”) and UiT Aurora project MASCOT. G. K. Verma was supported by UAEU grant G00004614. S. Ling is supported by Nanyang Technological University Research Grant No. 04INS000047C230GRT01 and UAEU grant G00004233.